LHC 2011 - latest news

Week 12: Machine coordinators: Mike Lamont – Ralph Assmann

Nine o'clock meetings on the weekend at nine o'clock.

LHC coordination vistar




Link to report from morning meeting

Monday 28th March

  • 06h10: Beams dumped for technical stop. Requested 1.38 TeV luminosity running completed.

8:30 meeting

Sunday 27th March

  • 10h02: Stable beams continued. vdM scans. Length calibrations.
  • 16h59: Beam dump. Delivered 21h 39min of 1.38TeV physics in this fill. Spontaneous LBDS trigger.
  • 18h15: Transfer line reference measurements before technical stop.
  • 19h45: Injection for next 1.38 TeV fill.
  • 21h16: Stable beams. 1.38 TeV.

9:00 meeting

Saturday 26th March

  • 10h35: Power converter RB.A78 intervention completed.
  • 16h15:  Next 1.38 TeV attempt
  • 18h20:  Stable beams. Fill #1653.

9:00 meeting

Friday 25th March

  • 01h30: Collimation completed and qualified.
  • 02h00: Beam dump tests.
  • 06h10: Stable beams at 1.38 TeV.
  • 09h03: Beam dump. One bunch dropped too much in intensity and gave BPMS interlock (dump protection).
  • 10h36: Injection. Bunch length (too long, > 1.7ns) giving injection problems. SPS working. Improved but not fully solved. SIS interlocks because of timing mismatch of BPM’s (old acquisition stamp). Solved by SIS reboot.
  • 12h14: Abort gap cleaning made operational. Update BI-QP fixed display.
  • 14h11: Injection for physics.
  • 14h23: Trip of RB in 81. Water problem on bridge B (Thyristor bank). Access required.
  • 17h55: Injection.
  • 19h05: Ramp.
  • 20h03: Stable beams. Fill #1651.
  • 22h37: Beam dump: One bunch dropped too much in intensity and gave BPMS interlock (dump protection).
  • 23h31: Trip power converter RB.A78. Accesses and repair.

8:30 meeting

Thursday 24th March

  • 05h30: Parasitic studies on leveling by vertical beam-beam separation in IR8
  • 1647 dumped around 10:30
  • Setup for 1.38 TeV - couple of accesses required
  • Collimator setup starts late evening

8:30 meeting

Wednesday 23rd March

  • 06h10: Programmed dump.
  • 07h54: Injection. 200b.
  • 08h39: Ramp.
  • 09h18: Stable beams. 2.4e32 lumi.
  • 09h32: Beam dump due to RQTL11.R6B1 trip. EPC and CV intervention in PM65. Change of water manifold in PC rack. Rebooting BLM crates (missing/broken NFS mounts). New version of LSA with improvements for smoothing (K) in ramp and physics (Greg). Collimator dry rampof collimator functions for 1.38 TeV. PNO.a3 test on RQTL11.R2B2. QH power supply discharging on dipoles of S45 (C19L5) -> fixed with access. RF tests on cavity 4B2.
  • 18h35: Precycle.
  • 20h27: Injection.
  • 21h08: Incorporation problem.
  • 21h42: Ramp.
  • 22h20: Stable beams #1647. 2.56e32 lumi.

8:30 meeting

Tuesday 22nd March

  • 05h05: BPM controls problem
  • 11h00: Beam back
  • 12h10: Stable beams Fill 1644. 136 bunches. Initial luminosity   1.6e32.
  • 18h00: beams dumped
  • 19h12: Starting filling for 200 b
  • 21h15: Stable beams   fill 1645, 200 bunch, initial lumi  ~2.5 e32 (emittance ~2.4 micron)

8:30 meeting

Monday 21st March

  • 06h00: End of fill #1640: 3.7 pb-1.
  • 06h14: RB S12 and S23 tripped before loading rampdown table, active filter. MPE piquet. Fip_Com_Lost for both RBs... Piquet called again.
  • 09h18: Ramp down of LHCb and ALICE dipoles and compensators. Ramp down of ALICE solenoid.
  • 09h27: Recovered from RB trips. Prepare for injection while magnets ramp down.
  • 09h50: Problems in PS. Mains down in SPS.
  • 10h30: Total voltage INTLK reduced to 3.5 MV (P. Baudrenghien).
  • 10h33: Injection.
  • 11h49: Start ramp. Beams dumped from false collimator interlock.
  • 12h34: Release of the new FGC software to fix the RF trip problem of two days ago (S. Page). Deploy fix for low level collimator controls software (A. Masi).
  • 13h08: Problem with the ADT: the check for the injection settings is failing. W. Hoefle verifying.
  • 13h18: Injection 136b. Large oscillations at the injection of SPS (PS not stable): larger emittances (2.3-3.8 murad-m) with respect to nominal.
  • 14h38: Start ramp.
  • 15h13: Stable beams.
  • 15h13: LHC fast BCT phase tuning done (J.J. Gras).
  • 21h45: Beam dumped by EIC.
  • 21h49: Ramping up LHCb and ALICE magnets.
  • 22h39: Injection.
  • 23h19: SIS/controls problem (JMS)

8:30 meeting